# LtU-ILI: An All-in-One Framework for Implicit Inference in Astrophysics and Cosmology

### v1

@@ 2.4 Validation
We evaluate the constraints and coverage of our learned posteriors on a labeled test set of data-parameter pairs $\mathcal{D}_{\rm test}:=\{(\bm{x}_{i},\bm{\theta}_{i})\}_{i=1}^{N_{\rm test}}$. This test set is meant to represent a fully independent sampling of the real data distribution, unseen by the model prior to validation. This test set may originate from the same source as our labeled training set $\mathcal{D}_{\rm train}$ but must be held independent to avoid overfitting and underestimation of the uncertainty (see Huppenkothen et al.(2023)Huppenkothen, Ntampaka, Ho, Fouesneau, Nord, Peek, Walmsley, Wu, Avestruz, Buck et al. for a review of best practices). If the distribution of data-parameter pairs in the test set matches that of the predictions of our learned posterior, then our model is calibrated and can be reliably extended to new, unlabeled data.
@@
Lastly, given samples from a reference posterior, we can evaluate the discrepancy between a model’s prediction and its optimum. Reference posteriors can often be attained via long-run MCMC chains for explicit likelihood approaches, and then used to compare with ILI methods. Lueckmann et al.(2021)Lueckmann, Boelts, Greenberg, Goncalves, and Macke performed an extensive quantitative analysis on various sample-wise distances for benchmarking ILI methods and found that Classifier 2 Sample Tests (C2ST; Lopez-Paz and Oquab(2016)) were the most constraining measure of accuracy. C2ST is defined as the accuracy of a trained classifier (often a neural network itself) to distinguish between the true and inferred posterior samples. High C2ST suggests that the true and inferred posterior are very different, whereas a C2ST of $\sim 0.5$ suggests they are nearly indistinguishable. C2ST is later used to benchmark LtU-ILI in Section 4.1.
Posterior samples are also useful for quantifying the calibration of our model uncertainty in parameter space. We can construct a direct comparison of the predicted percentiles by our inference engine to the true error observed in our validation dataset. We first define the Probability Integral Transform (PIT; Cook et al.(2006)Cook, Gelman, and Rubin) as the cumulative density function (CDF) of our model posterior given an input $\bm{x}_{\rm o}$,
@@
where $\Theta$ is the Heaviside step function. Stated plainly, the PIT value counts the number of times that our posterior samples $\hat{\bm{\theta}}$ fall below the true parameter value $\bm{\theta}$. If our model posteriors are globally consistent with the truth, i.e. we match the true posterior everywhere in data-parameter space, then the distribution of PIT values evaluated on data-parameter pairs from our test set must be uniformly distributed (Zhao et al.(2021)Zhao, Dalmasso, Izbicki, and Lee):
@@
As a result of this property, the distribution of PIT values is a common goodness-of-fit metric for conditional density models (e.g. Cook et al.(2006)Cook, Gelman, and Rubin; Bordoloi et al.(2010)Bordoloi, Lilly, and Amara; Tanaka et al.(2018)Tanaka, Coupon, Hsieh, Mineo, Nishizawa, Speagle, Furusawa, Miyazaki, and Murayama). This is also known as Simulation-Based Calibration (Talts et al.(2018)Talts, Betancourt, Simpson, Vehtari, and Gelman).
To better interpret errors in the PIT distribution, we can use a percentile-percentile (P-P) plot, explicitly comparing the CDF of PIT values to the CDF of a uniform random variable. Intuitively, the P-P plot measures: ‘What percentile level is my posterior model assigning to the true value, and with what frequency does this occur in the test set?’ If the predictive posterior is well-calibrated, these should be the same, e.g., we should predict the true value below the 50-th percentile 50% of the time. If it is not, the shape of the P-P plot acts as a sensitive probe of global bias or over-/under-dispersion of predictive uncertainty (see Figure 1 of Zhao et al.(2021)Zhao, Dalmasso, Izbicki, and Lee for interpretations of error with P-P plots). This is an extremely practical tool for understanding and correcting biases and over/under-dispersion in complex ILI posteriors.
@@
It is also sufficient to use approximations to check multivariate posterior coverage. The Tests of Accuracy with Random Points (TARP; Lemos et al.(2023)Lemos, Coogan, Hezaveh, and Levasseur) method uses samples in high-dimensional parameter space to estimate expected coverage probabilities. By examining the density of posterior samples within regions of parameter space near true values, TARP constructs estimates of posterior coverage which are guaranteed to converge to the true posterior coverage with sufficient samples. With enough samples, this test is necessary and sufficient to guarantee an optimal posterior estimate.
@@ 3 Code Structure
The LtU-ILI software is a pipeline for training implicit inference models to infer scalar parameters from observational data. A typical inference pipeline consists of three main stages: Data, Inference, and Validation. The Data stage involves loading datasets from file or setting up on-the-fly simulators (e.g. for the methodologies in Section 2.3) in order to provide data-parameter pairs for implicit inference. The Inference stage then takes these data-parameter pairs, trains neural networks to connect them with amortized posterior or likelihood representations (as in Section 2.2), and saves the fitted models to file. Finally, the Validation stage loads a test dataset as well as the fitted models from Inference in order to evaluate goodness-of-fit metrics (e.g. constraint level, posterior coverage, etc.) and produces inference on unlabeled data. A schematic of this pipeline is shown in Figure 2.
Each stage is independent and adaptive, such that the configuration of one stage can be changed while keeping the others the same and the pipeline will still run. For example, if you had a fixed dataset but wanted to try out many different neural architectures, you would only need to change the Inference configuration. Similarly, if you wanted to see the gain in constraining power of using a short, conservative data vector versus a longer, more informative data vector, you could simply change the Data configuration to load the different datasets and the Inference and Validation would proceed exactly the same way. This design principle is extremely practical for doing apples-to-apples comparisons of the many design choices and hyperparameters associated with ILI, a key tenet of robust and responsible machine learning (Huppenkothen et al.(2023)Huppenkothen, Ntampaka, Ho, Fouesneau, Nord, Peek, Walmsley, Wu, Avestruz, Buck et al.).
The pipeline is designed to promote exploratory research while also supporting efficient production-level testing. We provide a fully functional Jupyter interface and detailed examples to access each stage of LtU-ILI. The Jupyter interface is a widely-used tool for examining new problems, building educational material, and facilitating collaborative efforts (Wang et al.(2019)Wang, Mittal, Brooks, and Oney). The Jupyter interface of LtU-ILI provides step-by-step guidance for establishing and testing a new inference pipeline that is familiar to novices and experts alike. This allows LtU-ILI to act both as a user-friendly toolkit to explore inference on a new problem as well as carry out thorough testing and hyperparameter tuning through its yaml-based configuration API. Users can generate many configurations of the pipeline and run multiple fittings in parallel on large-scale GPU compute clusters.
@@ 3.2 Inference
Once these are provided, the Inference stage knows how to pull data-parameter pairs from the Data stage, perform data preprocessing, construct each training algorithm, minimize losses (see Equations 4, 5, and 9), and produce a learned posterior model, $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$. Each inference algorithm automatically implements best practices for training, including data normalization (Singh and Singh(2020)), adaptive stochastic gradient descent (Kingma and Ba(2014)), and validation-guided early stopping (Bai et al.(2021)Bai, Yang, Han, Yang, Li, Mao, Niu, and Liu). The remainder of this section is dedicated to the core considerations of configuring the Inference stage.
@@
The Inference stage is built on three distinct computational backends: pydelfi (Alsing et al.(2019)Alsing, Charnock, Feeney, and Wandelt), sbi (Tejero-Cantero et al.(2020)Tejero-Cantero, Boelts, Deistler, Lueckmann, Durkan, Gonçalves, Greenberg, and Macke), and lampe (Rozet et al.(2021)Rozet, Delaunoy, Miller et al.). pydelfi is a Tensorflow package (Abadi et al.(2016)Abadi, Barham, Chen, Chen, Davis, Dean, Devin, Ghemawat, Irving, Isard et al.) for performing NLE using active learning and data compression techniques. sbi is a PyTorch package (Paszke et al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan, Killeen, Lin, Gimelshein, Antiga et al.) for performing NPE, NLE, and NRE, as well as their sequential analogs. lampe is also a PyTorch package focused on NPE methods, with substantial support for diverse embedding networks and normalizing flow models. The implementations of NLE in pydelfi and sbi are similar in theory (Papamakarios et al.(2019)Papamakarios, Sterratt, and Murray) but differ in applications of embedding networks. pydelfi has the option to train in a two-step process, wherein first it trains an embedding network to optimally compress data via Fisher information maximization (Charnock et al.(2018)Charnock, Lavaux, and Wandelt; Alsing and Wandelt(2018)) and then attaches it to the head of an NLE model, whereas sbi exclusively uses NLE embedding networks to compress parameters. pydelfi’s two-step compression-estimation process has been shown to significantly improve simulation-efficiency during sequential training for applications in cosmology (Alsing et al.(2018)Alsing, Wandelt, and Feeney). The lampe NPE in LtU-ILI is a custom implementation of Lueckmann et al.(2017)Lueckmann, Goncalves, Bassetto, Öcal, Nonnenmacher, and Macke, similar to SNPE-B of sbi but allowing for greater variety of configuration, embedding networks, and NDEs. LtU-ILI is a unifying framework for all of these mutually-exclusive backends, allowing, for the first time, convenient apples-to-apples comparison of each implementation of ILI.
@@
The Tensorflow and PyTorch backends have a variety of pre-implemented NDEs and ratio classifiers. The NDEs are either Mixture Density Networks (MDNs; See Section 2.2) or normalizing flows (See Papamakarios et al.(2021)Papamakarios, Nalisnick, Rezende, Mohamed, and Lakshminarayanan, for a review). The pydelfi architectures are implemented from scratch, whereas the sbi flow models utilize nflows (Durkan et al.(2020)Durkan, Bekasov, Murray, and Papamakarios) and lampe models utilize zuko (Rozet et al.(2022)Rozet et al.). Both backends also accept various standard and customizable prior distributions. The primary requirement of a custom prior is the ability to evaluate $\log p(\bm{\theta})$ (for NDE models) or a proportional proxy (for NLE/NRE models). If the assumed prior’s support is bounded (e.g., a top-hat prior), the code internally implements an affine parameter transformation to avoid assigning a nonzero probability to regions of parameter space beyond prior bounds.
@@
A key component of LtU-ILI is model ensembling, wherein we independently train multiple neural networks on the same dataset and combine their predictions to improve robustness to overfitting and model uncertainty. Singular NDEs are prone to overconfidence (Hermans et al.(2022)Hermans, Delaunoy, Rozet, Wehenkel, Begy, and Louppe), i.e., they have a tendency to underestimate the predictive uncertainty. Deep ensembling (Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell) is a well-tested schema for correcting overconfidence by averaging NDE predictions from multiple models to inflate error bars and correctly capture model uncertainty. It is an alternative to the computationally-expensive Bayesian neural networks (Blundell et al.(2015)Blundell, Cornebise, Kavukcuoglu, and Wierstra; Gal and Ghahramani(2016); Cobb and Jalaian(2021)), in which the full weight posterior is learned during training and sampled during inference. We recommend that users utilize deep ensembles to ensure robust inference within LtU-ILI.
@@
The adaptability of neural networks to various data inputs has greatly contributed to their growth in astronomy and cosmology. The convenient implementations of the various NDEs and classifiers in LtU-ILI can be further augmented with customizable embedding networks. These embedding networks can use classic Keras-like neural layers (Chollet et al.(2015)Chollet et al.) in both Tensorflow and PyTorch, greatly simplifying implementation. The integration with embedding networks in the lampe backend allow for exotic inputs, like graphs and sequences. In the code, we provide several examples of embedding networks, including connections to convolutional and graph neural networks (LeCun et al.(1998)LeCun, Bottou, Bengio, and Haffner; Kipf and Welling(2016)). We also note that all backends can be run on GPUs, to take advantage of computational speed-ups when using complex embedding networks.
@@ 3.3 Validation
The most impactful configuration of the Validation stage is the choice of sampling methods. LtU-ILI provides three classes of samplers: MCMC, variational inference (VI), and direct sampling. All models (NPE/NLE/NRE) in both backends are natively integrated with the emcee (Foreman-Mackey et al.(2013)Foreman-Mackey, Hogg, Lang, and Goodman) sampling package. emcee implements the Goodman and Weare(2010) MCMC sampler and improves efficiency through parallelization and affine transformations but does not take advantage of probability gradients. The sbi models also have access to PyTorch’s pyro samplers (Bingham et al.(2019)Bingham, Chen, Jankowiak, Obermeyer, Pradhan, Karaletsos, Singh, Szerlip, Horsfall, and Goodman), including slice samplers (Neal(2003)), Hamiltonian Monte Carlo (HMC; Neal et al.(2011)Neal et al.), and the No-U-Turn sampler (NUTS; Hoffman et al.(2014)Hoffman, Gelman et al.). All emcee and pyro samplers have a unified interface within LtU-ILI, wherein users can simply specify their choice of sampler as well as the number, length, and thinning of MCMC chains. Also, we provide the functionality for sbi models to fit and sample from NDEs through neural VI (Graves(2011)). VI is a fast, approximate alternative to MCMC sampling, particularly useful in the high-dimensional parameter regime. Lastly, solely for the NPE models in sbi, users can utilize direct sampling, which is faster than both MCMC and VI. We remark that the choice of sampling configuration might change according to the validation metric in question, so LtU-ILI allows for the flexibility to specify this on a case-by-case basis.

### v2

@@ 2.5 Validation
We evaluate the constraints and coverage of our learned posteriors on a labeled test set of data-parameter pairs $\mathcal{D}_{\rm test}:=\{(\bm{x}_{i},\bm{\theta}_{i})\}_{i=1}^{N_{\rm test}}$. This test set is meant to represent a fully independent sampling of the real data distribution, unseen by the model prior to validation. This test set may originate from the same source as our labeled training set $\mathcal{D}_{\rm train}$ but must be held independent to avoid overfitting and underestimation of the uncertainty (see Huppenkothen et al. 2023 for a review of best practices). If the distribution of data-parameter pairs in the test set matches that of the predictions of our learned posterior, then our model is calibrated and can be reliably extended to new, unlabeled data.
@@
Lastly, given samples from a reference posterior, we can evaluate the discrepancy between a model’s prediction and its optimum. Reference posteriors can often be attained via long-run MCMC chains for explicit likelihood approaches, and then used to compare with ILI methods. Lueckmann et al. (2021) performed an extensive quantitative analysis on various sample-wise distances for benchmarking ILI methods and found that Classifier 2 Sample Tests (C2ST; Lopez-Paz and Oquab, 2016) were the most constraining measure of accuracy. C2ST is defined as the accuracy of a trained classifier (often a neural network itself) to distinguish between the true and inferred posterior samples. High C2ST suggests that the true and inferred posterior are very different, whereas a C2ST of $\sim 0.5$ suggests they are nearly indistinguishable. C2ST is later used to benchmark LtU-ILI in Section 4.1.
Posterior samples are also useful for quantifying the calibration of our model uncertainty in parameter space. We can construct a direct comparison of the predicted percentiles by our inference engine to the true error observed in our validation dataset. We first define the Probability Integral Transform (PIT; Cook et al., 2006) as the cumulative density function (CDF) of our model posterior given an input $\bm{x}_{\rm o}$,
@@
where $\Theta$ is the Heaviside step function. Stated plainly, the PIT value counts the number of times that our posterior samples $\hat{\bm{\theta}}$ fall below the true parameter value $\bm{\theta}$. If our model posteriors are globally consistent with the truth, i.e. we match the true posterior everywhere in data-parameter space, then the distribution of PIT values evaluated on data-parameter pairs from our test set must be uniformly distributed (Zhao et al., 2021):
@@
As a result of this property, the distribution of PIT values is a common goodness-of-fit metric for conditional density models (e.g. Cook et al., 2006; Bordoloi et al., 2010; Tanaka et al., 2018). This is also known as Simulation-Based Calibration (Talts et al., 2018).
To better interpret errors in the PIT distribution, we can use a percentile-percentile (P-P) plot, explicitly comparing the CDF of PIT values to the CDF of a uniform random variable. Intuitively, the P-P plot measures: ‘What percentile level is my posterior model assigning to the true value, and with what frequency does this occur in the test set?’ If the predictive posterior is well-calibrated, these should be the same, e.g., we should predict the true value below the 50-th percentile 50% of the time. If it is not, the shape of the P-P plot acts as a sensitive probe of global bias or over-/under-dispersion of predictive uncertainty (see Figure 1 of Zhao et al. 2021 for interpretations of error with P-P plots). This is an extremely practical tool for understanding and correcting biases and over/under-dispersion in complex ILI posteriors.
@@
It is also sufficient to use approximations to check multivariate posterior coverage. The Tests of Accuracy with Random Points (TARP; Lemos et al., 2023) method uses samples in high-dimensional parameter space to estimate expected coverage probabilities. By examining the density of posterior samples within regions of parameter space near true values, TARP constructs estimates of posterior coverage which are guaranteed to converge to the true posterior coverage with sufficient samples. With enough test samples, the TARP method is necessary and sufficient to test the accuracy of posterior estimators.
@@ 3 Code Structure
The LtU-ILI software is a pipeline for training implicit inference models to infer scalar parameters from observational data. A typical inference pipeline consists of three main stages: Data, Inference, and Validation. The Data stage involves loading datasets from file or setting up on-the-fly simulators (e.g. for the methodologies in Section 2.4) in order to provide data-parameter pairs for implicit inference. The Inference stage then takes these data-parameter pairs, trains neural networks to connect them with posterior or likelihood representations (as in Section 2.2), and saves the fitted models to file. Finally, the Validation stage loads a test dataset as well as the fitted models from Inference in order to evaluate goodness-of-fit metrics (e.g. constraint level, posterior coverage, etc.) and produces inference on unlabeled data. A schematic of this pipeline is shown in Figure 2.
Each stage is independent and adaptive, such that the configuration of one stage can be changed while keeping the others the same and the pipeline will still run. For example, if you had a fixed dataset but wanted to try out many different neural architectures, you would only need to change the Inference configuration. Similarly, if you wanted to see the gain in constraining power of using a short, conservative data vector versus a longer, more informative data vector, you could simply change the Data configuration to load the different datasets and the Inference and Validation would proceed exactly the same way. This design principle is extremely practical for doing apples-to-apples comparisons of the many design choices and hyperparameters associated with ILI, a key tenet of robust and responsible machine learning (Huppenkothen et al., 2023).
The pipeline is designed to promote exploratory research while also supporting efficient production-level testing. We provide a fully functional Jupyter interface and detailed examples to access each stage of LtU-ILI. The Jupyter interface is a widely-used tool for examining new problems, building educational material, and facilitating collaborative efforts (Wang et al., 2019). The Jupyter interface of LtU-ILI provides step-by-step guidance for establishing and testing a new inference pipeline that is familiar to novices and experts alike. This allows LtU-ILI to act both as a user-friendly toolkit to explore inference on a new problem as well as carry out thorough testing and hyperparameter tuning through its yaml-based configuration API. Users can generate many configurations of the pipeline and run multiple fittings in parallel on large-scale GPU compute clusters.
@@ 3.2 Inference
Once these are provided, the Inference stage knows how to pull data-parameter pairs from the Data stage, perform data preprocessing, construct each training algorithm, minimize losses (see Equations 4, 5, and 9), and produce a learned posterior model, $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$. Each inference algorithm automatically implements best practices for training, including data normalization (Singh and Singh, 2020), adaptive stochastic gradient descent (Kingma and Ba, 2014), and validation-guided early stopping (Bai et al., 2021). The remainder of this section is dedicated to the core considerations of configuring the Inference stage.
@@
The Inference stage is built on three distinct computational backends: pydelfi (Alsing et al., 2019), sbi (Tejero-Cantero et al., 2020), and lampe (Rozet et al., 2021). pydelfi is a Tensorflow package (Abadi et al., 2016) for performing NLE using active learning and data compression techniques. sbi is a PyTorch package (Paszke et al., 2019) for performing NPE, NLE, and NRE, as well as their sequential analogs. lampe is also a PyTorch package focused on NPE methods, with substantial support for diverse embedding networks and normalizing flow models. The implementations of NLE in pydelfi and sbi are similar in theory (Papamakarios et al., 2019) but differ in applications of embedding networks. pydelfi has the option to train in a two-step process, wherein first it trains an embedding network to optimally compress data via Fisher information maximization (Charnock et al., 2018; Alsing and Wandelt, 2018) and then attaches it to the head of an NLE model, whereas sbi exclusively uses NLE embedding networks to compress parameters. pydelfi’s two-step compression-estimation process has been shown to significantly improve simulation-efficiency during sequential training for applications in cosmology (Alsing et al., 2018). The lampe NPE in LtU-ILI is a custom implementation of Lueckmann et al. (2017), similar to SNPE-B of sbi but allowing for greater variety of configuration, embedding networks, and NDEs. LtU-ILI is a unifying framework for all of these mutually-exclusive backends, allowing, for the first time, convenient apples-to-apples comparison of each implementation of ILI.
@@
The Tensorflow and PyTorch backends have a variety of pre-implemented NDEs and ratio classifiers. The NDEs are either Mixture Density Networks (MDNs; See Section 2.2) or normalizing flows (See Papamakarios et al., 2021, for a review). The pydelfi backend possesses its own neural architectures as originally implemented from scratch in Alsing et al. (2019), whereas the sbi and lampe models have software dependencies, respectively nflows (Durkan et al., 2020) and zuko (Rozet et al., 2022). Both backends also accept various standard and customizable prior distributions. The primary requirement of a custom prior is the ability to evaluate $\log p(\bm{\theta})$ (for NDE models) or a proportional proxy (for NLE/NRE models). If the assumed prior’s support is bounded (e.g., a top-hat prior), the code internally implements an affine parameter transformation to avoid assigning a nonzero probability to regions of parameter space beyond prior bounds.
@@
A key component of LtU-ILI is model ensembling, wherein we independently train multiple neural networks on the same dataset and combine their predictions to improve robustness to overfitting and model uncertainty. Singular NDEs are prone to overconfidence (Hermans et al., 2022), i.e., they have a tendency to underestimate the predictive uncertainty. Deep ensembling (Lakshminarayanan et al., 2017) is a well-tested schema for correcting overconfidence by averaging NDE predictions from multiple models to inflate error bars and correctly capture model uncertainty. It is an alternative to the computationally-expensive Bayesian neural networks (Blundell et al., 2015; Gal and Ghahramani, 2016; Cobb and Jalaian, 2021), in which the full weight posterior is learned during training and sampled during inference. We recommend that users utilize deep ensembles to ensure robust inference within LtU-ILI.
@@
The adaptability of neural networks to various data inputs has greatly contributed to their growth in astronomy and cosmology. The convenient implementations of the various NDEs and classifiers in LtU-ILI can be further augmented with customizable embedding networks. These embedding networks can use classic Keras-like neural layers (Chollet et al., 2015) in both Tensorflow and PyTorch, greatly simplifying implementation. The integration with embedding networks in the lampe backend allow for exotic inputs, like graphs and sequences. In the code, we provide several examples of embedding networks, including connections to convolutional and graph neural networks (LeCun et al., 1998; Kipf and Welling, 2016). We also note that all backends can be run on GPUs, to take advantage of computational speed-ups when using complex embedding networks.
@@ 3.3 Validation
The most impactful configuration of the Validation stage is the choice of sampling methods. LtU-ILI provides three classes of samplers: MCMC, variational inference (VI), and direct sampling. All models (NPE/NLE/NRE) in both backends are natively integrated with the emcee (Foreman-Mackey et al., 2013) sampling package. emcee implements the Goodman and Weare (2010) MCMC sampler and improves efficiency through parallelization and affine transformations but does not take advantage of probability gradients. The sbi models also have access to PyTorch’s pyro samplers (Bingham et al., 2019), including slice samplers (Neal, 2003), Hamiltonian Monte Carlo (HMC; Neal et al., 2011), and the No-U-Turn sampler (NUTS; Hoffman et al., 2014). All emcee and pyro samplers have a unified interface within LtU-ILI, wherein users can simply specify their choice of sampler as well as the number, length, and thinning of MCMC chains. Also, we provide the functionality for sbi models to fit and sample from NDEs through neural VI (Graves, 2011). VI is a fast, approximate alternative to MCMC sampling, particularly useful in the high-dimensional parameter regime. Lastly, solely for the NPE models in sbi, users can utilize direct sampling, which is faster than both MCMC and VI. We remark that the choice of sampling configuration might change according to the validation metric in question, so LtU-ILI allows for the flexibility to specify this on a case-by-case basis.
3.4 Choosing a Backend
As described above and later tested in Section 4, the algorithmic behavior of the sbi, pydelfi, and lampe backends are nearly identical for the same experimental configurations (i.e., the same training data, hyperparameters, neural architectures, etc.). However, each backend has unique strengths and weaknesses and may be more appropriate for different problems. For example, sbi is the most commonly-used backend, with a set of standardized training procedures and architectures which have been validated across many problems (Lueckmann et al., 2021). It also has custom samplers for NLE and NRE methods which are faster than the generic emcee samplers used in the pydelfi backend. For new users, sbi is a reliable choice for exploring new datasets or setting up standard ILI applications. In contrast, lampe’s routines have access to a broad scope of exotic NDEs and allow for much greater flexibility when designing training procedures and custom embedding architectures. In our experience, using the flow models unique to lampe often produced tighter and better-calibrated posterior estimates relative to similar sbi applications. In addition, graph and recurrent embedding networks are currently only possible through the lampe interface, because of its capacity to handle complex data modalities. As a result, lampe is recommended for more advanced use of ILI in complex datasets. Lastly, our pydelfi tests in Section 4.2 appear to indicate stronger performance in sequential learning tasks than similar approaches with sbi. Also, pydelfi’s Tensorflow implementation may make it a more attractive option to users already working in Tensorflow, rather than the other options in PyTorch. Despite these rules of thumb, LtU-ILI is a fully unified interface incorporating all these backends. Our code allows users to easily explore the strengths of each backend for their specific applications.